OpenStudy (anonymous):

Solve each system of equations by substitution y = –2x – 5 y = 3x + 20

OpenStudy (anonymous):

You have \(y\) equals two different expressions, that means the two expression must have the same value. That's \(3x + 20=–2x – 5 \implies 5x=-25 \implies x=-5\). Plug this value of x in either equation, you get \(y=-2(-5)-5=5\). So \(x=-5\) and \(y=5\).